Consider a hydrogen atom in the 2p state. Determine at what radius will the electron probability be a maximum.
1
Expert's answer
2020-05-27T10:33:37-0400

Radial wave function of hydrogen atom in 2p state is "R_{21}=\frac{1}{\sqrt{3}}(\frac{1}{2a})^{\frac{3}{2}}(\frac{r}{a})e^{-(\frac{r}{2a})}"

The radial probability is

"|R_{21}|^2r^2=\frac{1}{{3}}(\frac{1}{2a})^{3}(\frac{r}{a})^2e^{-(\frac{r}{a})}r^2"

To find the maximum value of above equation, Differentiating above equation with respect to r and equating to zero, we get

"\frac{d}{dr}|R_{21}|^2r^2=0"

"\frac{d}{dr}[\frac{1}{{3}}(\frac{1}{2a})^{3}(\frac{r}{a})^2e^{-(\frac{r}{a})}r^2]=0"

"\frac{d}{dr}[r^4e^{-(\frac{r}{a})}]=0"

"[4r^3e^{-(\frac{r}{a})}+r^4(-\frac{1}{a})e^{-(\frac{r}{a})}]=0"

"4r^3+r^4(-\frac{1}{a})=0 \implies r-4a=0"

"r=4a" since "a" is the Bohr radius.

Hence, An electron has maximum probability at radius is "r=4a" .